OpenStudy (anonymous):

The heat of combustion is 22.7kJ/g which of the following is the theoretical mass of water that could be heated from 25 degrees to 53 degrees using the heat produced by the complete combustion of 0.5g of methanol a) 97g b)0.097g c)194g d)0.194g

OpenStudy (anonymous):

First, let's determine the amount of thermal energy released by combustion 0.5g of methanol. \[Q = m \cdot Q_{HV}\]where \(m\) is the mass of the methanol in grams and \(Q_{HV}\) is the heat of combustion of methanol as given. Second, let's come up with an expression that relates heat input into a sample of water to it's associated temperature change. \[Q_{i n} = m_{water} \cdot c_{water} \cdot (T_f -T_i) \] Finally, realize that all the heat released during combustion flows into the water, we can see that\[m_{methanol} \cdot Q_{HV} = m_{water} \cdot c_{water} \cdot (T_f - T_i)\] I'll leave it to you to solve for the mass of the water.

OpenStudy (anonymous):

First, you should understand everything eashmore said. Then use the eqation Q = mC(deltaT) If you rearrange for mass you get Q/C(change in temp) = m. Solving you get \[(22.7KJ/g \times .5g)\div(( 4.184J/g C)\times (53C-25C)) = 97g\]